What is the general form of the equation: -1/2(y+2)=(x-4)^2?

Answer by checkley71(8403)

-1/2(Y+2)=(X-4)^2
-(Y+2)/2=X^2-8X+16
(-Y-2)/2=X^2-8X+16 CROSS MULTIPLY
-Y-2=2X^2-16X+32
-Y=2X^2-16X+32+2
Y=-2X^2+16X-34 ANSWER.